A painting is currently worth $100,000, and is expected to maintain its real value for three years.
The real interest rate is expected to remain constant at 10%. What is the present value of the painting's expected price at the end of the third year?
Answer: $88,899

CAN YOU EXPLAIN HOW TO DO IT STEP BY STEP.
HOW DID YOU GET THE ANSWER 88,899????
PLEASE EXPLAIN IN DETAIL
Answer #1

Since the painting is expected to maintain its real value for three years

The value at the end of three years(Future value) = $ 100,000

Given, Real Interest rate = 10%

Thus, Present Value of the painting = 100,000/(1+ interest rate)^ 3 = 100,000/(1.1)^3 = $ 75131.48

The answer is not $ 88,899. The steps are as above.